OpenStudy (anonymous):

The sum of 2 numbers is 18. The difference between 3 times the smaller number and the larger number is 10. Define variables and formulate a system of equations from this situation. What is the larger number? work please? :D

OpenStudy (s):

x + y = 18 3x - y = 10 sum the two equations: 4x = 28 x = 7 now plug in x = 7 back into the two equations : 7 + y = 18 3(7) - y = 10 you get y = 11 So the smaller number is 7, larger number is 11

OpenStudy (anonymous):

thanks